A city wants to double the area of a rectangular playground that is 72 feet by 48 feet by adding the same distance X to the length and the width. Write and solve an equation to find the value of X.
Answer by mananth(16946)

72*48 =3456
(72+x)(48+x)=2*3456
3456+72x+48x+x^2=6912
x^2+120x+3456=6912
x^2+120x-3456-0
x^2+144x-24x-3456=0
x(x+144)-24(x+144)=0
(x+144)(x-24)=0
x= 24 which is positive.